OpenStudy (anonymous):

car Q and z are put though an endurance test to see if they can travel at 120 km/h for 5.0 hours. Each car has a 45-L fuel tank. Car z must stop to refuel after traveling for 4.2 hours. Car Q, however, travels for 5.4 hours before running out of gas. For each car, calculate the average kilometers traveled for each liter of gas (km/L)

OpenStudy (anonymous):

Okay ill do car Q and you do car z. I need distance, d, in relation to speed, v, and time,t . Well if car Q travels at 120km per hour, for 1 hour, it travels 120km. If it travels at 120km for two hours, goes twice the distance, 240km. So generally the distance traveled is the speed times by the time. \[d = v \times t\] checking the units of each variable confirms this \[km = \frac{km}{h} \times h\] Okay so I want the distance traveled in 5.4 hours for car Q. \[d = v \times t = 120 km/h \times 5.4 h = 648 km\] cool, now I have the total distance I know that car Q went 648 km on 45 litres, \[45L = 648 km\] now To get 1 litre I need to devide 45L by 45... \[\frac{45L}{45} = 1L = \frac{648}{45}= 14.4 km/L\]